OpenStudy (shelby1290):

The height of a soccer ball kicked in the air is given by the quadratic equation h(t)=-4.9(t-2.1)^2+23, where time, t, is in seconds and height, h(t), is in metres Is the ball still in the air after 6 seconds?

OpenStudy (welshfella):

if you solve h(t) = 0 you'll find how long the ball is in the air

OpenStudy (welshfella):

what values of t will make h(t) = 0?

OpenStudy (welshfella):

h(t)=-4.9(t-2.1)^2+23 = 0 (t - 2.1)^2 = -23/-4.9 can you finish this?

OpenStudy (welshfella):

(t - 2.1)^2 = 4.694 t - 2.1 = sqrt 4.694

OpenStudy (shelby1290):

h(t)=-4.9(t-2.1)^2+23 = 0 (t - 2.1)^2 = -23/-4.9 (t+4.41)=4.69 @welshfella i don't know what to do next

OpenStudy (welshfella):

t - 4.41 is incorrect the square root of (t - 2.1)^2 is simply (t - 2.1)

OpenStudy (welshfella):

(t - 2.1)^2 means 'all of t - 2.1 is sqaured so the square root of this is simply t - 2.1

OpenStudy (welshfella):

ok?

OpenStudy (welshfella):

so t - 2.1 = sqrt 4.69 t = 2.17 + 2.1 = 4.27 seconds so the ball is not in the air after 6 seconds

OpenStudy (welshfella):

total time in the air = 4.27 seconds

OpenStudy (welshfella):

OK? any questions?

OpenStudy (shelby1290):

@welshfella oh okay I see, thank you!

OpenStudy (welshfella):

yw